OpenStudy (anonymous):

which of the following represents the zeros of f(x) = 6x^3 - 31x^2 + 4x + 5?

OpenStudy (anonymous):

@phi ?

OpenStudy (anonymous):

are there options that you're given?

OpenStudy (anonymous):

yes a. -5, 1/3, 1/2 b. 5, 1/3, 1/2 c. 5, 1/3, -1/2 d. 5, -1/3, 1/2

OpenStudy (anonymous):

In that case, you can just plug each value into the equation and see if it gives you zero ie) \[ a) \\ f(-5) = 6(-5)^3 - 31(-5)^2 + 4(-5) + 5 = ? \\ f(1/3) = 6(1/3)^3 - 31(1/3)^2 + 4(1/3) + 5=? \\ f(1/2) = 6(1/2)^3 - 31(1/2)^2 + 4(1/2) + 5=? \]

OpenStudy (anonymous):

a makes it equal 0?

OpenStudy (anonymous):

You have to do the arithmetic to find out ^_^

OpenStudy (anonymous):

hmmm i got b as the answer.

OpenStudy (anonymous):

but i think i mess up on 1/3 let me try it again

OpenStudy (anonymous):

yeah i got b

OpenStudy (anonymous):

I think b is incorrect - plugging in x=1/3 does not make that equal 0.

OpenStudy (anonymous):

i knew i messed up on 1/3. ok lemme try c

OpenStudy (anonymous):

well, off the bat, if b doesn't work, neither will c - they both have x=1/3 as a root, and 1/3 we found is not a root.

OpenStudy (anonymous):

so it has to be d then because its the only one without 1/3

OpenStudy (anonymous):

correct! ^_^

OpenStudy (anonymous):

yayyyy thanks man. could you help me with a couple more

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok one is which of the following is a polynomial with the roots 

OpenStudy (anonymous):

the choices are a. x^3 - 2x^2 - 3x + 6 b. x^3 - 3x^2 - 5x + 15 c. x^3 + 2x^2 - 3x - 6 d. x^3 + 3x^2 - 5x - 15

OpenStudy (anonymous):

So for this, it's pretty much the same process, you just have to plug in the roots into each of the equations.

OpenStudy (anonymous):

yeah, but im not sure how to do that with radicals :/

OpenStudy (anonymous):

what's \[(\sqrt 3 )^2=?\]

OpenStudy (anonymous):

81